Consider the following cash flows:
  

Year Cash Flow
0 –$ 28,600
1 15,100
2 13,800
3 10,200

    
What is the profitability index for the cash flows if the relevant discount rate is 11 percent? (Do not round intermediate calculations and round your answer to 3 decimal places, e.g., 32.161.)

  

Profitability index            

  

What is the profitability index if the discount rate is 16 percent? (Do not round intermediate calculations and round your answer to 3 decimal places, e.g., 32.161.)

   

Profitability index            

  

What is the profitability index if the discount rate is 23 percent? (Do not round intermediate calculations and round your answer to 3 decimal places, e.g., 32.161.)

  

Profitability index            

Please show how to get the calculations.
Answer #1

1: Profitability Index = PV of cash inflows/ Cash outflow

= (CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n) / Cash outflow

= (15100/1.11^1+ 13800/1.11^2 + 10200/1.11^3)/ 28600

= 32262.15/28600

=1.128

2: Profitability Index = PV of cash inflows/ Cash outflow

= (CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n) / Cash outflow

= (15100/1.16^1+ 13800/1.16^2 + 10200/1.16^3)/ 28600

=29807.60/28600

=1.042

3: Profitability Index = PV of cash inflows/ Cash outflow

= (CF1/(1+r)^1 + CF2/(1+r)^2 …………CFn/(1+r)^n) / Cash outflow

= (15100/1.23^1+ 13800/1.23^2 + 10200/1.23^3)/ 28600

=26879.29/28600

= 0.94